# From Stars to Molecules: AI Guided Device-Agnostic Super-Resolution Imaging

### v1

@@ Abstract
Super-resolution imaging has revolutionized the study of systems ranging from molecular structures to distant galaxies. However, existing deep-learning-based methods require extensive calibration and retraining for each imaging setup, limiting their practical deployment. We introduce a device-agnostic deep-learning framework for super-resolution imaging of point-like emitters that eliminates the need for calibration data or explicit knowledge of optical system parameters. Our model is trained on a diverse, numerically simulated dataset encompassing a broad range of imaging conditions, enabling robust generalization across different optical setups. Once trained, it reconstructs super-resolved images directly from a single resolution-limited camera frame with superior accuracy and computational efficiency compared to conventional methods. We experimentally validate our approach using a custom-built microscopy setup with ground truth emitter positions and demonstrate its versatility on astronomical and single-molecule localization microscopy datasets, achieving unprecedented resolution without prior information. Our findings establish a pathway toward universal, calibration-free super-resolution imaging, expanding its applicability across scientific disciplines.
I Introduction
Every optical system is affected by the fundamental constraint on its resolution, known as the diffraction limit. This limitation restricts the ability to resolve structures significantly smaller than the wavelength of light divided by the numerical aperture of the optical system. Advances in super-resolution techniques have garnered significant interest across various fields by enhancing resolution beyond this inherent limit [1, 2]. In biomedical imaging, these techniques improve the quality of medical scans and enable detailed visualizations of cellular structures [3, 4, 5]. In material science, they allow the study of nanoscale properties of materials [6]. In astronomy, super-resolution techniques enable detailed object observation from space telescopes and probes, as well as resolving individual stars in densely packed star clusters [7, 8]. Many more applications rely on imaging objects composed of point-like or single-emitter sources. Localization microscopy provides detailed visualization of cellular structures at the molecular level [9, 10, 11, 12]. In quantum physics, super-resolution aids in characterizing the optical properties and positioning of quantum dots [13] and enables precise site-resolved imaging of cold atoms in optical lattices [14, 15]. Whether the goal is to improve image resolution, localize emitting sources, or perform emitter tracking and counting, super-resolution techniques unlock new possibilities, providing unprecedented insight across diverse fields.
Super-resolution imaging can be achieved through various techniques. Interpolation-based approaches enhance resolution by inferring missing details from neighboring pixel information [16]. Tomographic methods reconstruct super-resolved images through inversion procedures of multiple perspectives [17, 18]. Statistical approaches leverage the stochastic nature of emitting sources [9, 19]. Numerous additional methods, including those based on deterministic algorithms [20], sparse representation [21], and deep learning [22, 23], have been developed. Regardless of the specific framework, super-resolution techniques can generally be classified into two categories: reconstruction and parameter estimation. Reconstruction aims to directly restore the whole super-resolved image [20, 24, 25], while estimation focuses on extracting key parameters and features, such as emitter positions [19, 26, 27]. In the following discussion, we will focus exclusively on the former, more universal approach: reconstruction techniques. As they do not require sparse or blinking sources, these techniques can cover a significantly broader range of applications.
@@ I Introduction
Achieving accurate reconstruction of an object based on a single intensity image requires substantial knowledge about the imaging system under consideration. Central to this knowledge is an accurate estimation of the point spread function, which characterizes the response of the system to a point source of light. The reconstruction quality is inherently linked to the accuracy of the point spread function estimate; any inaccuracies will lead to poor reconstruction results [28]. However, its precise estimation entails additional calibration measurements involving an isolated emitting point source. This condition can prove very demanding, especially in systems with high concentration of emitters or low signal-to-noise ratio, which often necessitates preparing an additional finely tuned specimen for calibration purposes. Additionally, imaging systems commonly exhibit variations in the shape of the point spread function across the image. Consequently, reconstructing larger areas may require separate calibration for each segment [29].
@@
Omitting the need for calibration data, some approaches can be adapted to perform a blind reconstruction, operating without explicit prior knowledge of the point spread function [30, 31]. Instead, these methods impose additional assumptions on the imaging system, typically in the form of a predefined point spread function shape. During reconstruction, parameters of the point spread function, such as its width or other defining characteristics, are iteratively estimated directly from the input image. As a result, these techniques introduce imperfections by simultaneously reconstructing the observed object and the point spread function of the imaging system. Altogether, blind reconstruction remains a non-convex optimization problem, posing a challenging task in real-life scenarios and often yielding inferior results compared to non-blind algorithms.
@@
Similar device-specific constraints arise with reconstruction using deep learning techniques. While these methods often outperform classical algorithms in both reconstruction and parameter estimation tasks [32, 22, 24, 33, 34, 35], their calibration presents an even more significant obstacle. Unlike traditional approaches, these techniques operate on a data-driven paradigm, learning the reconstruction mapping by extracting relevant information from observed samples. However, due to the device-specific nature of a typical training process, the model can only be applied to a particular point spread function profile and additional factors such as emitter power, background noise distribution, and concentration of emitters, each requiring prior estimation from the calibration datasets [36]. Modifying these parameters necessitates remeasuring the calibration set and retraining the model, making it a resource-expensive procedure.
This work addresses the challenges of super-resolution techniques requiring calibration measurements for accurate image reconstruction. We present a novel approach to developing deep learning models, illustrating that their device-specificity can be overcome. The proposed device-agnostic approach, which extensively benefits from numerically simulating training datasets, produces a model capable of accurate reconstructions without calibration data or prior knowledge of imaging system parameters. Additionally, the trained model exhibits inherent robustness to non-uniformity and instability of imaging components. We validate the resolution-enhancement ability of our approach using experimentally acquired images and benchmark its performance against the well-established Richardson-Lucy deconvolution method. Furthermore, we highlight the device-agnostic versatility of our model by applying it to publicly available astronomy and microscopy datasets, achieving ultra high-resolution reconstructions without calibration or parameter estimation.
II Results
@@
We developed a device-agnostic modeling network (DAMN) designed to reconstruct an intensity image of point-like emitting sources, as illustrated in Fig. 1. Unlike traditional approaches, the DAMN model operates using a single intensity frame without requiring additional information about the optical parameters of the imaging system. Without the need for calibration measurements or retraining, it can directly process images from diverse applications. Our model leverages the convolutional architecture of a neural network (refer to the Materials and Methods section), enabling it to reconstruct frames of varying shapes and sizes. Moreover, the reconstruction process is performed by a single-pass propagation through the neural network, resulting in rapid reconstruction of super-resolved images.
@@
Figure 1: Schematic representation of the DAMN approach and its application to intensity images of point-like emitting sources. (A) The model is trained using numerically simulated data pairs comprising resolution-limited noisy images alongside their super-resolved counterparts. Each training sample represents a unique combination of underlying optical parameters, such as the width of the point spread function and the signal-to-noise ratio. (B) Following training, this model is applied to enhance the resolution of experimental images acquired using a real-life imaging system.
@@
We train the DAMN model using pairs of targeted ideal objects and their corresponding resolution-limited images generated through numerical simulation. The ensemble of data pairs covers various optical parameters, such as emitter power and concentration, background noise distribution, and even different shapes and widths of the point spread function. Their ranges were selected to cover the majority of realistic situations (additional details in Materials and Methods section). The extent of the synthetic training data is unprecedented in the astronomy or emitter microscopy fields. Subsequently, we utilized incremental learning techniques [37] for the model to acquire knowledge about all possible parameter combinations. As a result of the optimization and training process, this single model accurately reconstructs super-resolved images independently of the underlying parameter values. Consequently, without the need for retraining or calibration data, the model can be applied to diverse imaging systems and is resistant to their parameter changes, including inhomogeneity across a field of view and instability, such as drifts and fluctuations in time.
@@
To characterize the effectiveness of the DAMN model, we conduct a comparative analysis with the Richardson-Lucy deconvolution based on Bayesian estimation with a uniform prior [38, 39]. This widely used iterative algorithm implements image reconstruction by repeatedly utilizing a known point spread function. We evaluate the results using a separate test set of image pairs not included in the training data to assess the generalization ability for new samples. We quantify the performance of both methods using the mean absolute error, calculated between the reconstructed image and its target object, averaged over the whole test dataset. In contrast to the Richardson-Lucy algorithm, which requires (and was supplied with) prior knowledge of the imaging system, our DAMN model operates independently of any device-specific information. Despite this distinction, the device-agnostic model (without the prior information) outperforms the classical Richardson-Lucy deconvolution (with complete information on the point spread function) in terms of reconstruction quality, as discussed in the following text. Additionally, due to the iterative nature of the Richardson-Lucy algorithm, it requires three orders of magnitude higher computational time on average, further emphasize the advantages of the DAMN approach, see the Materials and Methods section.
@@
Panels (A) - (C) in Fig. 2 illustrate the dependence of mean absolute error on the emitter power, the width of the point spread function, and the concentration of emitters, respectively. The results of these log-log graphs are characterized using an average value over the test dataset with a corresponding 90% confidence interval. Panel (A) contains dual horizontal axes representing the varying emitter power using both the signal-to-noise ratio (SNR) and the peak-to-noise ratio (PNR). For additional details regarding the explored parameters and evaluation of both approaches, see the Materials and Methods section. The DAMN model, represented by the red color, exhibits significantly better performance than the green-colored Richardson-Lucy algorithm. Similar results can also be observed in panels (B) and (C). Panel (B) depicts the error dependence on the point spread function width (consult the Materials and Methods section). The remaining panel (C) characterizes the dependence on the emitter concentration, referring to the collective number of emitters in a single image. As seen across all three panels, our device-agnostic approach provides up to two orders of magnitude more accurate reconstructions than the device-specific Richardson-Lucy deconvolution. The depicted errors were evaluated across the whole $50\times 50$ pixel image and can be, alternatively, expressed as per-pixel errors by dividing the values by the number of pixels.
@@
Figure 2: The dependence of the mean absolute error on the signal-to-noise ratio (SNR), the width of a Gaussian point spread function (PSF), the number of emitters in the image (concentration), and the continuous transition between an Airy and a Gaussian PSF, respectively. The resulting averages of the green Richardson-Lucy algorithm and the red DAMN model are accompanied by their 90% confidence intervals over the test set. Panel (A) is provided with a secondary horizontal axis recalculating the SNR values to the peak-to-noise ratio (PNR). Across all panels, the DAMN model consistently outperforms the Richardson-Lucy deconvolution by up to two orders of magnitude. The optical parameters not investigated in a given panel have the following values: SNR $=500$, the average noise intensity $=10$, the concentration $=50$, and the Gaussian PSF with $\sigma=2$ px.
@@
The panel (D) in Fig. 2 visualizes the error dependence on the point spread function shape, gradually transitioning from an Airy pattern to a Gaussian profile of the same width. As seen from this semi-logarithmic graph, the performance of both methods remains approximately constant during this transition. Such behavior is expected from the Richardson-Lucy deconvolution, as we always provide the correct point spread function profile. On the other hand, the DAMN model was trained using solely the exact Gaussian and Airy point spread functions constituting half of the dataset each. Despite this simplification in the training process, the graph unambiguously demonstrates the adaptive ability of the DAMN approach to generalize to previously unseen point spread function shapes. This adaptability is beneficial during training, as it significantly reduces the data and simulation complexity required for training device-agnostic models. Altogether, the presented panels display numerous benefits of the DAMN approach, proving it superior to the state-of-the-art algorithm.
@@
Demonstrating the validity of our approach beyond the simulated data, we applied it to reconstruct experimentally acquired images. Explicitly for this, we developed an advanced optical microscopy setup shown in Fig. 3, which provides complete control over the spatial distribution of emitting point-like sources. An incoherent light illuminates a digital micromirror device (DMD), onto which a mask representing the targeted ground truth image is imposed. By configuring the micromirrors, we can prepare a mask containing an arbitrary number of sources positioned at desired locations. This mask is then re-imaged by a high-demagnification and high-resolution preparation system, creating point-like emitters and thus forming a sample object. Subsequently, a low-resolution imaging microscope projects the emitters onto a camera. The resulting intensity profile captured by the camera serves as the resolution-limited input for reconstruction methods. Together with the ground truth mask, this setup provides experimental test data pairs, which allow for exact metric quantification of each reconstruction method performance. For details, see the Materials and Methods section.
@@
Using this setup, we investigated the performance of both methods while varying the concentration of emitters. Panel (A) in Fig. 4 illustrates the mean absolute error between reconstructed images and their corresponding target objects using the experimental data, shown as dot markers. For comparison, the continuous lines represent the results obtained using simulated data with corresponding optical parameters. The DAMN model outperforms the Richardson-Lucy deconvolution by orders of magnitude, even when applied to real-world images acquired from experimental measurements. This improvement is achieved despite operating significantly below the resolution limit of our optical system. To further emphasize these findings, panel (B) in Fig. 4 shows a representative experimental image containing nearly 200 emitters and its reconstruction by each method. As observed, the DAMN model offers a near-perfect reconstruction of the original image without the need for any calibration. The Richardson-Lucy deconvolution exhibits severe deviations and artifacts despite knowing the point spread function of the imaging setup. This experimental demonstration highlights the benefits of our device-agnostic approach, which significantly outperforms the established state-of-the-art algorithm requiring prior knowledge of the point spread function.
@@
Figure 4: (A) The red and green dots represent the mean absolute error between the DMD-imposed masks and the super-resolved images reconstructed by each method. These errors were evaluated across various emitter concentrations. The accompanying continuous lines depict error values derived from simulated data using optical parameters estimated for our imaging system. (B) A typical camera image containing nearly 200 emitters, alongside its corresponding DMD-imposed mask and each method reconstructions. The circled areas contain a magnified region for easier visual comparison. It is evident that the DAMN model significantly outperforms the Richardson-Lucy (RL) algorithm even in regions where mutual emitter distance is well-below the Rayleigh resolution limit $R=3.9$ px (inset scalebars).
@@
We have also explored the benefits of incorporating upsampling layers into the DAMN convolutional architecture, further extending its capabilities. The modified network reconstructs images with an eightfold increase in dimensions, leading to an even more prominent improvement in the resolution. We applied this model to an astronomy image of a dense star cluster from observations of the southern spiral galaxy NGC 300 in the Sculptor constellation by the European Southern Observatory [40], see Fig. 5 (A). Panel (B) depicts a magnified segment of the galaxy with marked regions of interest, while panel (C) presents their high-resolution reconstruction produced by the DAMN model. Despite the stars in all three regions being below the resolution limit of the imaging system, the DAMN reconstruction clearly reveals them. Notably, a detailed examination of region III shows two distinct intensity peaks. While the original data resemble a single structure, our reconstruction indicates that it might be two proximal stars hidden deep below the resolution limit. Confirming this hypothesis would require further analysis using additional data. While the DAMN model can process the entire $7600\times 7600$ pixel image, saturation artifacts in the central part of the original data (Fig. 5 (A)) led us to demonstrate the reconstruction of smaller outer regions of the galaxy. Moreover, as shown in (C), the resolution enhancement is so significant that proper visualization requires considerable zoom into the reconstructed areas.
@@
Figure 5: Demonstration of the DAMN model super-resolving capabilities on an astronomy image of a dense star cluster. (A) Intensity image of the spiral galaxy NGC 300 acquired by European Southern Observatory. (B) Magnified region of the galaxy, highlighting key areas of interest. (C) High-resolution reconstructions of these regions generated by the DAMN model, revealing details of individual stars. These reconstructions do not use any calibration data or prior information on the employed imaging system. The region III reconstruction indicates the presence of a secondary source completely beyond the optical resolution of the telescope.
@@
Additionally, we applied the same DAMN model, without retraining, to a publicly available dataset [41, 42] acquired for the single-molecule localization microscopy challenge. Fig. 6 depicts the performance evaluation using these 500 tubulin images of $128\times 128$ pixels with high molecule concentration (A). Panel (B) shows a reference image provided by the SOSplugin [43], a least-square localization method assuming a Gaussian point spread function, which participated in the challenge and is available as a plugin for ImageJ. Panel (C) depicts a tubulin reconstruction from Deep-STORM [24], a state-of-the-art deep learning model for image reconstruction in localization microscopy. Unlike our approach, Deep-STORM operates in a device-specific regime, requiring extensive optical system information, such as camera specification, point-spread-function model, approximate SNR, and expected emitter density. In comparison, the DAMN model directly processes the dataset without calibration or prior knowledge of the imaging system and provides the super-resolved tubulin image, see panel (D). The highlighted areas show a close-up comparison of all three approaches. Our DAMN model accurately captures the underlying tubulin structure, achieving resolution significantly surpassing that of the SOSplugin localization method. Unlike the localization method, the DAMN model clearly resolves the two vertically oriented, overlapping microtubules and reconstructs the horizontal one without gaps. Moreover, a close inspection reveals that our model provides a sharper image with a cleaner and thinner microtubule structure compared to the device-specific Deep-STORM. Panel (E) visualizes the projection of the microtubule profile over the yellow segment, demonstrating the resolution improvement. This result ultimately underlines the reconstruction ability of device-agnostic models in real-world applications. Although the presented model provides super-resolved reconstructions, the same paradigm of device-agnostic deep learning can be applied to localization tasks, i.e., directly predicting the positions and intensities of the emitting sources. Such a model could be highly advantageous for single-molecule localization microscopy of densely populated samples.
@@
Figure 6: Demonstration of the DAMN model on a high-concentration tubulin dataset from the single-molecule localization microscopy challenge. (A) A low-resolution image integrated from the set of 500 measured images, containing numerous emitting molecules. (B) A reference binary map of localized molecule positions derived from the dataset. (C) A reconstruction generated by Deep-STORM, a device-specific deep learning method calibrated using detailed information on the imaging setup. (D) A super-resolved image produced from the calibration-free, device-agnostic model without prior information. Panels (B-D) include inset images with a magnified region highlighting the resolution details. (E) Projection of the microtubule profile over the yellow rectangle segment.
III Conclusion

### v2

@@ Abstract
Super-resolution imaging has revolutionized the study of systems ranging from molecular structures to distant galaxies. However, existing super-resolution methods require extensive calibration and retraining for each imaging setup, limiting their practical deployment. We introduce a device-agnostic deep-learning framework for super-resolution imaging of point-like emitters that eliminates the need for calibration data or explicit knowledge of optical system parameters. Our model is trained on a diverse, numerically simulated dataset encompassing a broad range of imaging conditions, enabling generalization across different optical setups. Once trained, it reconstructs super-resolved images directly from a single resolution-limited camera frame with superior accuracy and computational efficiency compared to state-of-the-art methods. We experimentally validate our approach using a custom microscopy setup with ground-truth emitter positions. We also demonstrate its versatility on astronomical and single-molecule localization microscopy datasets, achieving unprecedented resolution without prior information. Our findings establish a pathway toward universal, calibration-free super-resolution imaging, expanding its applicability across scientific disciplines.
Main
Optical resolution is fundamentally constrained by diffraction. This limits the ability to observe structures comparable to the wavelength of light divided by the numerical aperture of the imaging system. Super-resolution imaging has emerged as a transformative technique in disciplines ranging from biomedical [1, 2] and material sciences [3] to astronomy [4, 5], by circumventing this limit and revealing previously inaccessible structural details. Numerous applications rely specifically on imaging point-like or single-emitter sources. In biology, single-molecule localization microscopy enables nanoscale visualization of cellular structures [6, 7, 8, 9, 10, 11]. Quantum physics leverages super-resolution for characterizing quantum dots [12] and precisely imaging cold atoms in optical lattices [13, 14]. Astronomy benefits through resolving individual stars and galaxies [15, 16].
Despite these successes, super-resolution methods require precise calibration and extensive knowledge of optical parameters. This requirement presents a substantial barrier, often involving laborious measurements, calibration data acquisition, and computationally expensive model retraining. Furthermore, practical applications frequently struggle with inhomogeneities and instabilities in imaging conditions, which further limit the versatility of existing algorithms.
To address these fundamental challenges, we propose a novel deep-learning framework that is inherently device-agnostic, eliminating the need for calibration or prior system-specific information. Our method exploits numerically simulated data encompassing an unprecedented range of optical conditions, enabling extreme generalization and adaptability across different imaging setups. Once trained, the model delivers rapid and accurate super-resolution reconstructions directly from a single, diffraction-limited image frame.
In this work, we demonstrate the efficacy of our approach through computational evaluations and experimental validations across multiple benchmarks, outperforming statistical Bayesian and deep-learning-based state-of-the-art methods. We further confirm the practical advantages and wide applicability of our approach using diverse datasets from molecular localization microscopy to astronomical imaging. The developed framework paves the way towards universal, calibration-free super-resolution imaging, significantly enhancing its accessibility and impact across scientific and technological disciplines.
Super-resolution and device-dependence
Super-resolution imaging can be achieved through various approaches, such as linear inverse and statistical Bayesian algorithms [17], sparse representation [18], tomographic image synthesis [19, 10, 20], and methods based on blinking emitters [6, 7, 8]. Many deep learning super-resolution methods have been developed in the last decade [21, 22, 23, 24, 25]. Recently, artificial intelligence has been used to identify novel super-resolution microscopy setups [26].
Regardless of the specific approach, super-resolution imaging can generally be classified into two categories: reconstruction and parameter estimation. Reconstruction aims to directly restore the whole super-resolved image [22, 17], while estimation focuses on extracting key parameters and features, such as emitter localization [27, 24, 28]. The localization performance rapidly decreases for images with higher number of overlapping emitters, which typically limits its application to stochastically blinking or photoactivated samples. In the following discussion, we will focus exclusively on the former, more universal approach: image reconstruction.
@@
Achieving accurate reconstruction of an object based on a single intensity image requires substantial knowledge about the imaging system under consideration. Central to this knowledge is an accurate estimation of the point spread function, which characterizes the response of the system to a point source of light. The reconstruction quality is inherently linked to the accuracy of the point spread function estimate; any inaccuracies will lead to poor reconstruction results [29]. However, its precise estimation entails additional calibration measurements involving an isolated emitting point source. This condition can prove very demanding, especially in systems with high concentration of emitters, low signal-to-noise ratio, or temporally developing systems. Additionally, imaging systems commonly exhibit variations in the shape of the point spread function across the image. Consequently, reconstructing larger areas may require separate calibration for each segment [30].
@@
Some approaches can be adapted to perform a blind reconstruction, operating without explicit prior knowledge of the point spread function [31]. Instead, these methods impose additional assumptions on the imaging system, typically in the form of a predefined point spread function shape. During reconstruction, parameters of the point spread function, such as its width or other defining characteristics, are iteratively estimated directly from the input image. As a result, these methods introduce imperfections by simultaneously reconstructing the observed object and the point spread function of the imaging system. Altogether, blind reconstruction remains a non-convex optimization problem, posing a challenging task in real-life scenarios and often yielding inferior results compared to non-blind algorithms.
@@
Similar device-dependent constraints arise with reconstruction using deep learning. While these approaches often outperform classical algorithms in both reconstruction and parameter estimation tasks [21, 22, 24, 25], their calibration presents an even more significant obstacle. Unlike traditional approaches, deep learning model operates on a data-driven paradigm, inferring the reconstruction mapping by extracting relevant information from observed samples. However, due to the device-dependent nature of a typical training process, the model can only be applied to a particular point spread function profile and additional conditions such as emitter power, background noise distribution, and concentration of emitters, each requiring prior estimation from the calibration datasets [32]. Modifying these parameters requires remeasuring the calibration set and retraining the model, making it a resource-expensive procedure.
Results
Device-agnostic modeling
@@ Results
We developed a device-agnostic modeling network (DAMN) designed to reconstruct an intensity image of point-like emitting sources, as illustrated in Fig. 1. Unlike traditional approaches, the DAMN model operates using a single intensity frame without requiring additional information about the optical parameters of the imaging system. Without the need for calibration measurements or retraining, it can directly process images from diverse applications. Our model leverages the convolutional architecture of a neural network (refer to the Methods section), enabling it to reconstruct frames of varying shapes and sizes. Moreover, the reconstruction process is performed by a single-pass propagation through the neural network, resulting in rapid reconstruction of super-resolved images.
@@
Figure 1: Schematic representation of the DAMN approach and its application to intensity images of point-like emitting sources. a The model is trained using numerically simulated data pairs comprising resolution-limited noisy images alongside their super-resolved counterparts. Each training sample represents a unique combination of underlying optical parameters, such as the width of the point spread function and the signal-to-noise ratio. b Following training, this model is applied to enhance the resolution of experimental images acquired using a real-life imaging system.
@@
We train the DAMN model using pairs of targeted ideal objects and their corresponding resolution-limited images generated through numerical simulation. The ensemble of data pairs covers various optical parameters, such as emitter power and concentration, background noise distribution, and even different shapes and widths of the point spread function. Their ranges were selected to cover the majority of realistic situations (additional details in the Methods section). The extent of the synthetic training data is unprecedented in the astronomy or emitter microscopy fields. Subsequently, we utilized incremental learning techniques [33] for the model to acquire knowledge about the widest possible parameter combinations. As a result of the optimization and training process, this single model accurately reconstructs super-resolved images independently of the underlying parameter values. Consequently, without the need for retraining or calibration data, the model can be applied to diverse imaging systems and is resistant to their parameter changes, including inhomogeneity across a field of view and instability, such as drifts and fluctuations in time.
Evaluation using simulated datasets
@@
To characterize the effectiveness of the DAMN model, we conduct a comparative analysis with the Richardson-Lucy deconvolution based on Bayesian estimation with a uniform prior [34, 35]. This widely used iterative algorithm implements image reconstruction by repeatedly utilizing a known point spread function. We evaluate the results using a separate test set of image pairs not included in the training data to assess the generalization ability for new samples. We quantify the performance of both methods using the mean absolute error, calculated between the reconstructed image and its target object, averaged over the whole test dataset. In contrast to the Richardson-Lucy algorithm, which requires (and was supplied with) prior knowledge of the imaging system, our DAMN model operates independently of any device-dependent information. Despite this distinction, the device-agnostic model (without the prior information) outperforms the classical Richardson-Lucy deconvolution (with complete information on the point spread function) in terms of reconstruction quality, as discussed in the following text. Additionally, due to the iterative nature of the Richardson-Lucy algorithm, it requires three orders of magnitude higher computational time on average, further emphasize the advantages of the DAMN approach, see the Methods section.
@@
Panels a - c in Fig. 2 illustrate the dependence of mean absolute error on the emitter power, the width of the point spread function, and the concentration of emitters, respectively. The results of these log-log graphs are characterized using an average value over the test dataset with a corresponding 90% confidence interval. Panel a contains dual horizontal axes representing the varying emitter power using both the signal-to-noise ratio (SNR) and the peak-to-noise ratio (PNR). For additional details regarding the explored parameters and evaluation of both approaches, see the Methods section. The DAMN model, represented by the red color, exhibits significantly better performance than the green-colored Richardson-Lucy algorithm. Similar results can also be observed in panels b and c. Panel b depicts the error dependence on the point spread function width (consult the Methods section). The remaining panel c characterizes the dependence on the emitter concentration, referring to the collective number of emitters in a single image. As seen across all three panels, our device-agnostic approach provides up to two orders of magnitude more accurate reconstructions than the device-dependent Richardson-Lucy deconvolution. The depicted errors were evaluated across the whole $50\times 50$ pixel image and can be, alternatively, expressed as per-pixel errors by dividing the values by the number of pixels.
@@
Figure 2: The dependence of the mean absolute error on a the signal-to-noise ratio (SNR), b the width of a Gaussian point spread function (PSF), c the number of emitters in the image (concentration), and d the continuous transition between an Airy and a Gaussian PSF, respectively. The resulting averages of the Richardson-Lucy algorithm (green) and the DAMN model (red) are accompanied by their 90% confidence intervals over the test set. Panel a is provided with a secondary horizontal axis recalculating the SNR values to the peak-to-noise ratio (PNR). Across all panels, the DAMN model consistently outperforms the Richardson-Lucy deconvolution by up to two orders of magnitude. The optical parameters not investigated in a given panel have the following values: SNR $=500$, the average noise intensity $=10$, the concentration $=50$, and the Gaussian PSF with $\sigma=2$ px.
@@
The panel d in Fig. 2 visualizes the error dependence on the point spread function shape, gradually transitioning from an Airy pattern to a Gaussian profile of the same width. As seen from this semi-logarithmic graph, the performance of both methods remains approximately constant during this transition. Such behavior is expected from the Richardson-Lucy deconvolution, as we always provide the correct point spread function profile. On the other hand, the DAMN model was trained using solely the exact Gaussian and Airy point spread functions constituting half of the dataset each. Despite this simplification in the training process, the graph unambiguously demonstrates the adaptive ability of the DAMN approach to generalize to previously unseen point spread function shapes. This adaptability is beneficial during training, as it significantly reduces the data and simulation complexity required for training device-agnostic models. Altogether, the presented panels display numerous benefits of the DAMN approach, proving it superior to the state-of-the-art algorithm.
Optical microscopy experimental validation
@@
Demonstrating the validity of our approach beyond the simulated data, we applied it to reconstruct experimentally acquired images. Explicitly for this, we developed a custom-built optical microscopy setup illustrated in Fig. 3 and shown in detail in Extended data Fig. 1. This microscope provides complete control over the spatial distribution of emitting point-like sources, achieving this for the first time in super-resolution benchmarking. An incoherent light illuminates a digital micromirror device (DMD), onto which a mask representing the targeted ground-truth image is imposed. By configuring the micromirrors, we prepare a mask containing an arbitrary number of sources positioned at desired locations. This mask is then re-imaged by a high-demagnification and high-resolution preparation system, creating point-like emitters and thus forming a sample object. Subsequently, a low-resolution imaging microscope projects the emitters onto a camera. The resulting intensity profile captured by the camera serves as the resolution-limited input for reconstruction methods. Together with the ground-truth mask, this setup provides experimental test data pairs, which allow for exact metric quantification of each reconstruction method performance. For details, see the Methods section.
@@
Using this setup, we investigated the performance of both methods while varying the concentration of emitters. Panel a in Fig. 4 illustrates the mean absolute error between reconstructed images and their corresponding target objects using the experimental data, shown as dot markers. For comparison, the continuous lines represent the results obtained using simulated data with corresponding optical parameters. The DAMN model outperforms the Richardson-Lucy deconvolution by orders of magnitude, even when applied to real-world images acquired from experimental measurements. This improvement is achieved despite operating significantly below the resolution limit of our optical system. To further emphasize these findings, panel b in Fig. 4 shows a representative experimental image containing nearly 200 emitters and its reconstruction by each method. As observed, the DAMN model offers a near-perfect reconstruction of the original image without the need for any calibration. The Richardson-Lucy deconvolution exhibits severe deviations and artifacts despite knowing the point spread function of the imaging setup. This experimental demonstration highlights the benefits of our device-agnostic approach, which significantly outperforms the established state-of-the-art algorithm requiring prior knowledge of the point spread function.
@@
Figure 4: a The red and green dots represent the mean absolute error between the DMD-imposed masks and the super-resolved images reconstructed by each method. These errors were evaluated across various emitter concentrations. The accompanying continuous lines depict error values derived from simulated data using optical parameters estimated for our imaging system. b A typical camera image containing nearly 200 emitters, alongside its corresponding DMD-imposed mask and each method reconstructions. The circled areas contain a magnified region for easier visual comparison. It is evident that the DAMN model significantly outperforms the Richardson-Lucy (RL) algorithm even in regions where mutual emitter distance is well-below the Rayleigh resolution limit $R=3.9$ px (inset scalebars).
Astronomical and localization microscopy demonstration
@@
We have also explored the benefits of incorporating upsampling layers into the DAMN convolutional architecture, further extending its capabilities. The modified network reconstructs images with an eightfold increase in dimensions, leading to an even more prominent improvement in the resolution. We applied this model to an astronomy image of a dense star cluster from observations of the southern spiral galaxy NGC 300 in the Sculptor constellation by the European Southern Observatory [36], see Fig. 5 a. Panel b depicts a magnified segment of the galaxy with marked regions of interest, while panel c presents their high-resolution reconstruction produced by the DAMN model. Despite the stars in all three regions being below the resolution limit of the imaging system, the DAMN reconstruction clearly reveals them. Notably, a detailed examination of region III shows two distinct intensity peaks. While the original data resemble a single structure, our reconstruction indicates that it might be two proximal stars, with an approximate 0.11 arcsec angular distance, hidden deep below the resolution limit. Confirming this hypothesis would require further analysis using additional data. While the DAMN model can process the entire $7600\times 7600$ pixel image, saturation artifacts in the central part of the original data (Fig. 5 a) led us to demonstrate the reconstruction of smaller outer regions of the galaxy. Moreover, as shown in c, the resolution enhancement is so significant that proper visualization requires considerable zoom into the reconstructed areas.
@@
Figure 5: Demonstration of the DAMN model super-resolving capabilities on an astronomy image of a dense star cluster. a Intensity image of the spiral galaxy NGC 300 acquired by European Southern Observatory. b Magnified region of the galaxy, highlighting key areas of interest. c High-resolution reconstructions of these regions generated by the DAMN model, revealing details of individual stars. These reconstructions do not use any calibration data or prior information on the employed imaging system. The region III reconstruction indicates the presence of a secondary source completely beyond the optical resolution of the telescope.
@@
Additionally, we applied the same DAMN model, without retraining, to a publicly available dataset [37, 38] acquired for the single-molecule localization microscopy challenge. Fig. 6 depicts the performance evaluation using these 500 tubulin images of $128\times 128$ pixels with high molecule concentration a. Panel b shows a reference image provided by the SOSplugin [39], a least-square localization method assuming a Gaussian point spread function, which participated in the challenge and is available as a plugin for ImageJ. Panel c depicts a tubulin reconstruction from Deep-STORM [22], a state-of-the-art deep learning model for image reconstruction in localization microscopy. Unlike our approach, Deep-STORM operates in a device-dependent regime, requiring extensive optical system information, such as camera specification, point-spread-function model, approximate SNR, and expected emitter density. In comparison, the DAMN model directly processes the dataset without calibration or prior knowledge of the imaging system and provides the super-resolved tubulin image, see panel d. The highlighted areas show a close-up comparison of all three approaches. Our DAMN model accurately captures the underlying tubulin structure, achieving resolution significantly surpassing that of the SOSplugin localization method. Unlike the localization method, the DAMN model clearly resolves the two vertically oriented, overlapping microtubules and reconstructs the horizontal one without gaps. Moreover, a close inspection reveals that our model provides a sharper image with a cleaner and thinner microtubule structure compared to the device-dependent Deep-STORM. Panel e visualizes the projection of the microtubule profile over the yellow segment, demonstrating the resolution improvement. This result ultimately underlines the reconstruction ability of device-agnostic models in real-world applications. Although the presented model provides super-resolved reconstructions, the same paradigm of device-agnostic deep learning can be applied to localization tasks, i.e., directly predicting the positions and intensities of the emitting sources. Such a model could be highly advantageous for single-molecule localization microscopy of densely populated samples.
@@
Figure 6: Demonstration of the DAMN model on a high-concentration tubulin dataset from the single-molecule localization microscopy challenge. a A low-resolution image integrated from the set of 500 measured images, containing numerous emitting molecules. b A reference binary map of localized molecule positions derived from the dataset. c A reconstruction generated by Deep-STORM, a device-dependent deep learning approach calibrated using detailed information on the imaging setup. d A super-resolved image produced from the calibration-free, device-agnostic model without prior information. Panels b-d include inset images with a magnified region highlighting the resolution details. e Projection of the microtubule profile over the yellow rectangle segment.
Discussion